OpenStudy (anime_lover_levi):

(2sqrt(5)+3sqrt(7))^2 Simplify!!!! HELP WILL MEDAL

OpenStudy (deathwillwaitforyou):

17?

OpenStudy (anime_lover_levi):

no.... @DeathWillWaitForYou

OpenStudy (lorenbeech):

What simple form do you need it in? an approximate answer or a square root answer?

OpenStudy (anime_lover_levi):

square root ,l ike a problem @lorenbeech

OpenStudy (lorenbeech):

2sqrt(5)+21

OpenStudy (anime_lover_levi):

No , i no that , idk how to get there though

OpenStudy (anime_lover_levi):

...

jhonyy9 (jhonyy9):

use (a+b)^2 = a^2 +2ab +b^2 (2sqrt(5)+3sqrt(7))^2 = (2sqrt5)^2 +2*2sqrt5 *3sqrt7 +(3sqrt7)^2 = = 4*5 +12sqrt35 +9*7 = =20 +12sqrt35 +63 = = 83 +12sqrt35

OpenStudy (anime_lover_levi):

thank you so much

jhonyy9 (jhonyy9):

np anytime was my pleasure good luck bye bye